### v2

@@ Abstract
Healthcare systems worldwide face persistent challenges in efficiency, accessibility, and personalization. Modern artificial intelligence (AI) has shown promise in addressing these issues through precise predictive modeling; however, its impact remains constrained by limited integration into clinical workflows. Powered by modern AI technologies such as multimodal large language models and world models, Embodied AI (EmAI) represents a transformative frontier, offering enhanced autonomy and the ability to interact with the physical world to address these challenges. As an interdisciplinary and rapidly evolving research domain, “EmAI in healthcare” spans diverse fields such as algorithms, robotics, and biomedicine. This complexity underscores the importance of timely reviews and analyses to track advancements, address challenges, and foster cross-disciplinary collaboration. In this paper, we provide a comprehensive overview of the “brain” of EmAI for healthcare, wherein we introduce foundational AI algorithms for perception, actuation, planning, and memory, and focus on presenting the healthcare applications spanning clinical interventions, daily care & companionship, infrastructure support, and biomedical research. These significant advancements have the potential to enable personalized care, enhance diagnostic accuracy, and optimize treatment outcomes. Despite its promise, the development of EmAI for healthcare is hindered by critical challenges such as safety concerns, gaps between simulation platforms and real-world applications, the absence of standardized benchmarks, and uneven progress across interdisciplinary domains. We discuss the technical barriers and explore ethical considerations, offering a forward-looking perspective on the future of EmAI in healthcare. A hierarchical framework of intelligent levels for EmAI systems is also introduced to guide further development. By providing systematic insights, this work aims to inspire innovation and practical applications, paving the way for a new era of intelligent, patient-centered healthcare.
@@ 1 Introduction
Healthcare services play a fundamental role in human well-being, yet they face persistent challenges, including inequities in access [1], inefficiencies in care delivery [2], and a growing demand for personalized solutions to address complex medical conditions [3, 4]. These issues primarily stem from limited and unevenly distributed healthcare resources [5], as well as insufficiently advanced treatment methods [6], often resulting in delayed, inadequate, or sometimes excessive treatments that exacerbate patients’ conditions [7]. Within the confines of current clinical workflows—largely reliant on finite clinical infrastructure, human healthcare professionals, and caregiving staff—these challenges remain difficult to fully overcome. To address these issues, various efforts have been implemented, such as telemedicine services [8, 9], automated triage systems [10, 11], AI-assisted healthcare monitoring [12], and medical image analysis [13, 14], which have enhanced the precision and efficiency of medical access while contributing to the ongoing transformation of the healthcare landscape. However, they still fall short of providing direct support within existing clinical workflows.
@@
Artificial intelligence (AI) technologies, particularly deep learning approaches, are introducing a new workforce into healthcare practice, driving the ongoing transformation of the healthcare landscape [15, 16, 17, 18, 19, 20, 21]. These methods learn medical and diagnostic knowledge from extensive healthcare data collected across multiple centers, devices, scenarios, patients, and time points, utilizing electronic health records (EHRs), genomic sequences, health monitoring signals, and medical images to perform advanced clinical predictive modeling [22, 23]. This enables early-stage diagnoses [24], facilitates personalized treatment recommendations [25], identifies subtle disease manifestations beyond human discernment [26], and advances biomedical research [27], collectively improving both the efficiency and quality of healthcare services.
@@
However, the translation of modern AI technologies into tangible clinical benefits remains constrained by at least four fundamental challenges: (I) Insufficient multimodal processing. Current AI systems primarily rely on one or several common modalities such as vision, language, and audio, but often lack the capability to process tactile sensations and olfactory cues, which are both more complex and critical in healthcare. The absence of integration for these less-explored modalities limits the effectiveness of AI in addressing the multifaceted nature of clinical tasks and patient care. (II) The separation between development and deployment. Current deep learning frameworks are characterized by a clear separation between development and inference phases, which hinders their continuous evolution in real-world clinical settings. This rigid separation delays adaptation to dynamic clinical requirements and evolving environments, ultimately restricting the systems’ capacity for ongoing self-improvement. (III) Insufficient human-machine interaction functionalities. Effective interaction with patients and healthcare professionals is critical for enhancing patient experiences and even improving treatment outcomes. While cutting-edge conversational AI systems, such as ChatGPT and GPT-4, exhibit remarkable interaction capabilities, they often fall short in aligning with treatment objectives and extending beyond verbal communication to encompass behavioral interactions. Such interactions demand advanced reasoning, robust memory retention, and the ability to adapt based on experience. Although recent studies have highlighted the transformative potential of language in therapeutic contexts [28], the mechanisms through which an AI system’s linguistic and interactive behaviors can positively impact clinical outcomes—particularly in areas like mental health treatment—remain underexplored [29, 30]. (IV) The absence of pathways from decision-making to action execution. Without embodiment in robotic or assistive devices, AI systems are unable to directly alleviate the workload of healthcare professionals and caregiving staff. While current deep learning models may provide accurate diagnoses and decision support, they rarely translate these insights into actionable diagnostic or therapeutic interventions. Furthermore, ensuring safety during such interventions and maintaining seamless integration within established clinical workflows remain critical challenges that require urgent resolution.
@@
Embodied AI (EmAI) is emerging as a promising approach to addressing these challenges in healthcare scenarios [31, 32, 33, 34, 35, 36]. By integrating AI algorithms, especially multimodal large language models (MLLMs) and world models, with innovations from robotics, mechatronics, human-computer interaction, and sensor technologies, EmAI equips AI algorithms with a physical “body” or tangible medium, enabling direct interaction with the world [37]. The AI algorithms are responsible for executing perception, action control, decision-making, and memory processing, ensuring the seamless operation of EmAI systems. Several recent breakthroughs in AI algorithms have significantly advanced the development of EmAI. For instance, unsupervised learning has enabled AI to extract foundational knowledge from vast data without human supervision [38, 39, 40, 41, 42]; interactive perceptual learning [43] has empowered EmAI systems to comprehend causal relationships of objects and assess the interaction possibilities and feasibility of engaging with various objects [44]; cross-modality fusion techniques have been extensively developed to integrate and leverage complementary information from diverse sources [45, 46]; deep reinforcement learning allows AI systems to learn optimal behaviors through feedback from the environment [47, 48, 49, 50, 51, 52]; and advancements in large language models (LLMs) [53, 54, 55, 56, 57], MLLMs [58, 59, 60, 61, 62, 63], vision-language-action (VLA) models [64, 65, 66, 67] and even world models [68, 69, 70] have provided AI systems with enhanced communication capabilities, reasoning abilities, and action-planning capabilities, particularly for tasks like navigation and manipulation [71, 72]. Thanks to these achievements, the development and usability of the “EmAI brain” have significantly advanced, enabling more sophisticated, adaptive, and context-aware EmAI systems  [^1] In this review, we primarily focus on the core component of an EmAI system—the AI system, often referred to as the “brain” of EmAI. For the sake of clarity, we do not make a strict distinction between the “brain” of EmAI and the overall “EmAI” system in the following discussion. capable of functioning in dynamic healthcare environments.
@@
Figure 1: Overview of embodied AI in healthcare research. (a) The proportion of healthcare-specific EmAI studies among all EmAI publications (as indexed by Google Scholar) from 2019 to 2024, reflecting the growing focus on healthcare within this field. (b) Publication trends (2019–2024) for EmAI applications in healthcare. The growing research interest is illustrated across four key areas in healthcare: Biomedical Research, Infrastructure Support, Daily Care & Companionship, and Clinical Intervention, reflecting the increasing recognition of EmAI’s potential to address diverse healthcare challenges. (c) A keyword co-occurrence network was generated using VOSviewer [73], based on Web of Science data spanning 2019 to 2024, showcasing core concepts and applications of EmAI in healthcare. Keywords extracted from article titles and abstracts are represented as nodes, with node size indicating frequency and links representing co-occurrence relationships. The network demonstrates a strong trend of interdisciplinary collaboration among fields such as computer science, engineering, and robotics to build holistic EmAI systems for the healthcare applications.
@@
Advances in EmAI are driving transformative applications across various fields, with healthcare emerging as a leading domain, which accounts for approximately 35% of the field’s work [74, 27], as illustrated in Figure 1(a). Notable examples include surgical robots [75] and companion robots [76], which are becoming increasingly widespread. Figure 1(b) highlights the remarkable growth of EmAI research in key healthcare domains, including biomedical research, infrastructure support, daily care & companionship, and clinical intervention. Notably, the total number of publications in 2024 is nearly sevenfold that of 2019, with clinical intervention research showing the fastest growth while maintaining a substantial share across these domains. Such achievements are underpinned by the integration of insights from multiple disciplines. As depicted in the keyword co-occurrence network (Figure 1(c)), the dense interconnections across domains highlight how advancements in one field catalyze progress in others, emphasizing the pivotal role of interdisciplinary collaboration in revolutionizing healthcare. Notable contributions stem from breakthroughs in foundation models, large language models, computer vision, cognitive science, sociology, and robotics, collectively shaping the future of EmAI applications in healthcare. Building on these research achievements, EmAI has been profoundly transforming healthcare by enhancing patient care and operational efficiency. It enables robotic diagnostics [77], precise surgical interventions [78], and personalized rehabilitation therapies [79], not only streamlining medical workflows but also delivering improved health outcomes and reduced recovery times [80]. Beyond clinical applications, EmAI provides meaningful companionship [81] and emotional support [82], offering particular benefits to vulnerable groups such as children, the elderly, and individuals with disabilities or chronic illnesses, thereby alleviating the burden on healthcare providers. In addition, EmAI is redefining biomedical research by automating experimental processes and analyzing large-scale datasets, enabling researchers to generate insights and conduct experiments with unprecedented speed. These advancements have accelerated the discovery of medical mechanisms [83, 84, 85], therapeutic targets [86, 87], and disease prevention strategies [88, 89], driving innovation across the biomedical landscape.
@@
Despite significant advancements [35, 90, 91, 92, 93, 94, 95, 96], the development of EmAI for healthcare remains in its infancy and faces multiple challenges. Current efforts often concentrate on isolated components of EmAI [80, 97, 98], such as developing advanced algorithms [99, 100], improving workflows [101, 102], or curating data sets [103, 104], without achieving integration into comprehensive systems. To realize the full potential of EmAI, cross-disciplinary collaboration is essential to bridge these fragmented contributions and build cohesive, end-to-end solutions. Moreover, research has predominantly focused on high-profile applications, such as surgical robotics [77, 105, 106, 78, 107, 108, 109, 110], while other promising areas, including mental health interventions [91, 92], remain underexplored. This uneven distribution of attention limits the broader impact of EmAI across diverse healthcare needs. Additionally, while companion robots have shown potential, most are reactive rather than proactive [81], restricting their ability to anticipate and address patient needs autonomously [111, 112]. Similarly, biomedical research robots face difficulties in maintaining precision and reliability in the inherently complex and dynamic environments of medical research.
@@
Additionally, the development of EmAI for healthcare continues to encounter significant technical challenges. First, EmAI development is typically carried out on simulation platforms, which often fail to accurately replicate real-world environments. This discrepancy presents a major challenge in bridging the gap between simulations and real-world applications. Additionally, as EmAI systems may directly interact with the real world, ensuring safety becomes even more critical, especially in medical tasks [113, 114, 115]. Second, although EmAI systems rely on large datasets, the acquisition of large, ethically sourced, domain-specific real-world datasets in healthcare is hindered by privacy regulations and complex clinical workflows, creating significant barriers to the development of healthcare-specific EmAI. Other challenges, such as ethical considerations [116, 117, 118, 119] and economic and societal implications [120, 121], were also expected to be addressed.
@@
Given the promising potential and numerous benefits of EmAI for both patients and healthcare professionals, as well as the existing challenges, a timely summary of these aspects is crucial for advancing the field and fostering interdisciplinary collaboration. In this review, we summarize and discuss recent and emerging applications of EmAI in healthcare, highlighting key factors that could significantly impact patient outcomes and healthcare practices. In Section 2, we provide a concise overview of the technologies that underpin the “EmAI brain”, covering four essential capabilities: perception, actuation, high-level planning, and memory. While our focus is not on delving into the technical foundations of EmAI (for technical reviews, see [122, 123, 65]) or its general applications in robotics (refer to [124, 125, 37]), we present the first comprehensive review centered on healthcare applications of modern EmAI, particularly in clinical interventions, daily care & companionship, infrastructure support, and biomedical research (discussed in Section 3). We also summarize their progress and limitations, categorizing EmAI into five levels of intelligence and illustrating each with examples from various healthcare domains (see Section 4). This framework aims to guide researchers and practitioners in understanding the evolution and stages of EmAI in healthcare. Datasets and benchmarks for diverse healthcare scenarios are summarized in Section 5, while the challenges and opportunities are further outlined in Section 6, with the aim of guiding researchers towards relevant fields, applications, and data foundations for future exploration.

### v3

@@ Abstract
Healthcare systems worldwide face persistent challenges in efficiency, accessibility, and personalization. Modern artificial intelligence (AI) has shown promise in addressing these issues through precise predictive modeling; however, its impact remains constrained by limited integration into clinical workflows. Powered by modern AI technologies such as multimodal large language models and world models, Embodied AI (EmAI) represents a transformative frontier, offering enhanced autonomy and the ability to interact with the physical world to address these challenges. As an interdisciplinary and rapidly evolving research domain, “EmAI in healthcare” spans diverse fields such as algorithms, robotics, and biomedicine. This complexity underscores the importance of timely reviews and analyses to track advancements, address challenges, and foster cross-disciplinary collaboration. In this paper, we provide a comprehensive overview of the “brain” of EmAI for healthcare, wherein we introduce foundational AI algorithms for perception, actuation, planning, and memory, and focus on presenting the healthcare applications spanning clinical interventions, daily care & companionship, infrastructure support, and biomedical research, covering 35 specialized tasks. These significant advancements have the potential to enable personalized care, enhance diagnostic accuracy, and optimize treatment outcomes. Despite its promise, the development of EmAI for healthcare is hindered by critical challenges such as safety concerns, gaps between simulation platforms and real-world applications, the absence of standardized benchmarks, and uneven progress across interdisciplinary domains. We discuss the technical barriers and explore ethical considerations, offering a forward-looking perspective on the future of EmAI in healthcare. A hierarchical framework of intelligent levels for EmAI systems is also introduced to guide further development. By providing systematic insights, this work aims to inspire innovation and practical applications, paving the way for a new era of intelligent, patient-centered healthcare.
@@ 1 Introduction
Healthcare services play a fundamental role in human well-being, yet they face persistent challenges, including inequities in access [1], inefficiencies in care delivery [2], and a growing demand for personalized solutions to address complex medical conditions [3, 4]. These issues primarily stem from limited and unevenly distributed healthcare resources [5], as well as insufficiently advanced treatment methods [6], often resulting in delayed, inadequate, or sometimes excessive treatments that exacerbate patients’ conditions [7]. Within the confines of current clinical workflows—largely reliant on finite clinical infrastructure, human healthcare professionals, and caregiving staff—these challenges remain difficult to fully overcome. To address these issues, various efforts have been implemented, such as telemedicine services [8, 9], automated triage systems [10, 11], AI-assisted healthcare monitoring [12], and medical image analysis [13, 14], which have enhanced the precision and efficiency of medical access. However, they still fall short of providing direct support within existing clinical workflows.
@@
Artificial intelligence (AI) technologies, particularly deep learning approaches, are introducing a new workforce into healthcare practice, driving the ongoing transformation of the healthcare landscape [15, 16, 17, 18, 19, 20, 21]. These methods learn medical and diagnostic knowledge from extensive healthcare data collected across multiple centers, scenarios, devices, and time points, utilizing electronic health records (EHRs), genomic sequences, health monitoring signals, and medical images to perform advanced clinical predictive modeling [22, 23]. This enables early-stage diagnoses [24], supports personalized treatment recommendations [25], and detects subtle disease manifestations beyond human detection [26], collectively improving the quality of healthcare services.
@@
However, the translation of modern AI technologies into tangible clinical benefits remains constrained by at least four fundamental challenges: (I) Insufficient multimodal processing. Current AI systems primarily rely on one or several common modalities such as vision, language, and audio, but often struggle to process tactile sensations and olfactory cues, which are both more complex and critical in healthcare. The insufficient integration for these less-explored modalities limits the effectiveness of AI in addressing the multifaceted nature of clinical tasks and patient care. (II) The separation between development and deployment. Current deep learning frameworks are characterized by a distinct divide between development and inference phases, which hinders their continuous adaptation in real-world clinical settings. This rigid separation delays adaptation to dynamic clinical requirements and evolving environments, ultimately restricting the systems’ capacity for ongoing self-improvement. (III) Insufficient human-machine interaction functionalities. Effective interaction with patients and healthcare professionals is critical for enhancing patient experiences and even improving treatment outcomes. While cutting-edge conversational AI systems, such as ChatGPT and GPT-4, exhibit remarkable interaction capabilities, they often fall short in aligning with clinical goals and extending beyond verbal communication to encompass behavioral interactions. Effective interactions require advanced reasoning, robust memory retention, and the ability to adapt based on experience. Although recent studies have highlighted the transformative potential of language in therapeutic contexts [27], the mechanisms through which an AI system’s linguistic and interactive behaviors can positively impact clinical outcomes—particularly in areas like mental health treatment—remain underexplored [28, 29]. (IV) The absence of pathways from decision-making to action execution. Without embodiment in robotic or assistive devices, AI systems are unable to directly alleviate the workload of healthcare professionals and caregiving staff. While current deep learning models may provide accurate diagnoses and decision support, they rarely translate these insights into actionable diagnostic or therapeutic interventions. Furthermore, ensuring safety during such interventions and maintaining seamless integration within established clinical workflows remain critical challenges that require urgent resolution.
@@
Embodied AI (EmAI) is emerging as a promising approach to addressing these challenges in healthcare scenarios [30, 31, 32, 33, 34, 35]. By integrating AI algorithms, especially multimodal large language models (MLLMs) and world models, with innovations from robotics, mechatronics, human-computer interaction, and sensor technologies, EmAI equips AI algorithms with a physical “body”, enabling direct interaction with the world [36]. The AI algorithms are responsible for executing perception, action control, decision-making, and memory processing, ensuring the seamless operation of EmAI systems. Several recent breakthroughs in AI algorithms have significantly advanced the development of EmAI. For instance, unsupervised learning has enabled AI to extract foundational knowledge from vast data without human supervision [37, 38, 39, 40, 41]; interactive perceptual learning [42] has empowered EmAI systems to comprehend causal relationships of objects and assess the interaction possibilities and feasibility of engaging with various objects [43]; cross-modality fusion techniques have been extensively developed to integrate and leverage complementary information from diverse sources [44, 45]; deep reinforcement learning allows AI systems to learn optimal behaviors through feedback from the environment [46, 47, 48, 49, 50, 51]; and advancements in large language models (LLMs) [52, 53, 54, 55, 56], MLLMs [57, 58, 59, 60, 61, 62], vision-language-action (VLA) models [63, 64, 65, 66] and world models [67, 68, 69] have provided AI systems with enhanced reasoning, action-planning, and communication capabilities, particularly for tasks like navigation and manipulation [70, 71]. Thanks to these achievements, the development and usability of the “EmAI brain” have significantly advanced, enabling more sophisticated, adaptive, and context-aware EmAI systems  [^1] In this review, we primarily focus on the core component of an EmAI system—the AI system, often referred to as the “brain” of EmAI. For the sake of clarity, we do not make a strict distinction between the “brain” of EmAI and the overall “EmAI” system in the following discussion. capable of functioning in dynamic healthcare environments.
@@
Figure 1: Overview of embodied AI in healthcare research. Top left: The proportion of healthcare-specific EmAI studies among all EmAI publications (as indexed by Google Scholar) from 2019 to 2024, reflecting the growing focus on healthcare within this field. Bottom left: Publication trends (2019–2024) for EmAI applications in healthcare. The growing research interest is illustrated across four key areas in healthcare: Biomedical Research, Infrastructure Support, Daily Care & Companionship, and Clinical Intervention, reflecting the increasing recognition of EmAI’s potential to address diverse healthcare challenges. Right: A keyword co-occurrence network was generated using VOSviewer [72], based on Web of Science data spanning 2019 to 2024, showcasing core concepts and applications of EmAI in healthcare. Keywords extracted from article titles and abstracts are represented as nodes, with node size indicating frequency and links representing co-occurrence relationships. The network demonstrates a strong trend of interdisciplinary collaboration among fields such as computer science, engineering, and robotics to build holistic EmAI systems for the healthcare applications.
@@
Advances in EmAI are driving transformative applications across various fields, with healthcare emerging as a leading domain, which accounts for approximately 35% of the field’s work [73, 74], as illustrated in Figure 1(a). Notable examples include surgical robots [75] and companion robots [76], which are becoming increasingly widespread. Figure 1(b) highlights the remarkable growth of EmAI research in key healthcare domains, including biomedical research, infrastructure support, daily care & companionship, and clinical intervention. Notably, the total number of publications in 2024 is nearly sevenfold that of 2019, with clinical intervention research showing the fastest growth while maintaining a substantial share across these domains. Such achievements are built on the integration of insights from multiple disciplines. As depicted in the keyword co-occurrence network (Figure 1(c)), the dense interconnections across domains reveal how advancements in one field catalyze progress in others, underscoring the pivotal role of interdisciplinary collaboration in revolutionizing healthcare. Notable contributions emerge from breakthroughs in foundation models, large language models, computer vision, cognitive science, sociology, and robotics, collectively shaping the future of EmAI applications in healthcare. Building on these research achievements, EmAI has been significantly transforming healthcare by enhancing patient care and operational efficiency. It enables robotic diagnostics [77], precise surgical interventions [78], and personalized rehabilitation therapies [79], which not only streamline medical workflows but also enhance health outcomes and reduce recovery times [80]. Beyond clinical applications, EmAI provides meaningful companionship [81] and emotional support [82], which is especially beneficial to vulnerable groups such as children, the elderly, and individuals with disabilities or chronic illnesses, thus helping to ease the burden on healthcare providers. In addition, EmAI is transforming biomedical research by automating experimental processes and analyzing large-scale datasets, which allows researchers to generate insights and conduct experiments with unprecedented speed. These advancements have accelerated the discovery of medical mechanisms [83, 84, 85], therapeutic targets [86, 87], and disease prevention strategies [88, 89], driving innovation across the biomedical landscape.
@@
Despite significant advancements [34, 90, 91, 92, 93, 94, 95, 96], the development of EmAI for healthcare remains in its infancy and faces multiple challenges. Current efforts often concentrate on isolated components of EmAI [80, 97, 98], such as developing advanced algorithms [99, 100], improving workflows [101, 102], or curating data sets [103, 104], without achieving integration into comprehensive systems. To realize the full potential of EmAI, cross-disciplinary collaboration is essential to bridge these fragmented contributions and build cohesive, end-to-end solutions. Moreover, research has largely focused on high-profile applications, such as surgical robotics [77, 105, 106, 78, 107, 108, 109, 110], while other promising areas, such as mental health interventions [91, 92], remain underexplored. This uneven distribution of attention limits the broader impact of EmAI across diverse healthcare needs. Additionally, while companion robots have shown potential, most are reactive rather than proactive [81], which restricts their ability to anticipate and address patient needs autonomously [111, 112]. Similarly, biomedical research robots face challenges in maintaining precision and reliability in the inherently complex and dynamic environments of medical research.
@@
Additionally, the development of EmAI for healthcare still faces significant technical challenges. First, EmAI development is typically carried out on simulation platforms, which often fail to replicate real-world environments accurately. This discrepancy poses a major challenge in bridging the gap between simulations and real-world applications. Additionally, as EmAI systems may directly interact with the real world, ensuring safety becomes even more critical, especially in medical tasks [113, 114, 115]. Second, although EmAI systems rely on large datasets, the acquisition of large, ethically sourced, domain-specific real-world datasets in healthcare is obstructed by privacy regulations and complex clinical workflows, creating significant barriers to the development of healthcare-specific EmAI. Other challenges, such as ethical considerations [116, 117, 118, 119] and economic and societal implications [120, 121],must also be addressed.
@@
Given the promising potential and numerous benefits of EmAI for both patients and healthcare professionals, as well as the existing challenges, a timely summary of these aspects is crucial for advancing the field and fostering interdisciplinary collaboration. In this review, we summarize and discuss recent and emerging applications of EmAI in healthcare, emphasizing key factors that could significantly impact patient outcomes and healthcare practices. In Section 2, we provide a concise overview of the technologies that underpin the “EmAI brain”, covering four essential capabilities: perception, actuation, high-level planning, and memory. While our focus is not on exploring the technical foundations of EmAI (for technical reviews, see [122, 123, 64]) or its general applications in robotics (refer to [124, 125, 36]), we present a comprehensive review centered on healthcare applications of modern EmAI, particularly in clinical interventions, daily care & companionship, infrastructure support, and biomedical research (discussed in Section 3). We also summarize their progress and limitations, categorizing EmAI into five levels of intelligence and illustrating each with examples from various healthcare domains (see Section 4). This framework aims to guide researchers and practitioners in understanding the evolution and stages of EmAI in healthcare. Datasets and benchmarks for diverse healthcare scenarios are summarized in Section 5, while the challenges and opportunities are further outlined in Section 6, with the aim of guiding researchers towards relevant fields, applications, and data foundations for future exploration.
Highlights. This survey is the first to comprehensively integrate technological, application-driven, and ethical perspectives on Embodied AI in healthcare. Specifically, we:
  • 1.
    Techniques: Provide an in-depth review of foundational algorithms for multimodal perception, actuation, planning, and memory, which form the core of EmAI’s intelligent systems.
  • 2.
    Applications: Examine how EmAI is revolutionizing healthcare across four critical domains, covering 35 specialized tasks and underscoring its transformative value in diverse scenarios.
  • 3.
    Datasets and Benchmarks: Summarize and categorize key datasets and benchmarks to facilitate and accelerate research in diverse application fields.
  • 4.
    Evaluation Framework: Propose a novel framework for assessing the intelligence levels of EmAI systems, guiding future development.
  • 5.
    Challenges and Opportunities: Discuss technical barriers, safety concerns, and ethical considerations, offering a forward-looking perspective on the future of EmAI in healthcare.